OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1).

OpenStudy (callisto):

Equation: \[\frac{y-y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1}\] Put (x1, y1) = (4, 1) and (x2, y2) = (5, -1) \[\frac{y-1}{x-4} = \frac{-1-1}{5-4}\] Can you continue from here?

OpenStudy (anonymous):

deffiently not ...

OpenStudy (callisto):

Hmm.. Do you understand the above steps?

OpenStudy (anonymous):

y = 4x + 3

OpenStudy (callisto):

Where does y = 4x + 3 come from? Can you show your workings?

OpenStudy (callisto):

Or please tell me which part you don't understand, so that I could explain it..

OpenStudy (anonymous):

I was doing the wrong question on that last response . Lol Sorry.

OpenStudy (anonymous):

Lol i don't understand any of it.. Im taking Higher classes than what im suppose to take ... So I just google all of my test and homework awnsers so i don't fail.. I do online ..

OpenStudy (callisto):

Then... perhaps you should learn it...

OpenStudy (callisto):

Basically, you need to find a slope first. Slope = (y2 - y1)/(x2 - x1) When (x1, y1) is the coordinates for a point. (x2, y2) is the coordinates for another point

OpenStudy (callisto):

Now in your case, let (x1, y1) = (4, 1) and (x2, y2) = (5, -1) Can you put the values into the corresponding places? Show your workings please.

OpenStudy (savvy):

slope of a line is (y1-y2)/(x1-x2) for a line which passes through (x1,y1) and (x2,y2)... slope int. form says y = slope(x) + y int. slope is -2 now y = -2x + y int. now put y =1 and x =4 since the line passes thru (4,1) and hence the eqn. satisfies it....putting the values we get y int. = 9 so eqn. is y = -2x + 9